OpenStudy (anonymous):

Solve x2 − 8x − 20 = 0. x = 4, x = −5 x = −4, x = 5 x = 2, x = −10 x = −2, x = 10

OpenStudy (anonymous):

@mathmate @Teddyiswatshecallsme

OpenStudy (welshfella):

you need to find 2 numbers whose product = -20 and whose sum = -8

OpenStudy (welshfella):

these will be the values of a and b in the factors of the expression (x + a)(x + b) = 0 you can then find the zeroes

OpenStudy (welshfella):

try factoring -10 and see if you can make -8 from 2 of the factors

OpenStudy (welshfella):

sorry thats -20 not -10

OpenStudy (anonymous):

Ok I think I understand what you mean...... ok so -10 * 2 = -20....... and then I have to find a way to make -20 equal -8....?

OpenStudy (welshfella):

you are on the right track -10 and 2 is right because -10 + 2 = -8

OpenStudy (welshfella):

so that the factors are (x - 10 )(x + 2)

OpenStudy (anonymous):

Oh wow that was suprisingly simple.....

OpenStudy (welshfella):

yea when you expand the factors the -10*x and 2*x make -8x

OpenStudy (anonymous):

Ah ok thanks man

OpenStudy (welshfella):

yw now to get the zeros you equate to 0 (x - 10)(x + 2) = 0 means that x-10 = 0 or x + 2 = 0 so the zeros are 10 and -2